# Fast Scrambling in Classically Simulable Quantum Circuits

### v1

@@ II Super-Clifford circuits
In the context of our super-Clifford circuits, the familiar notion of a stabiliser state is replaced by a ‘stabiliser operator’. Specifically, for an $N$ qubit system, we define a stabiliser operator as an an operator in $\mathcal{S}$ which is fixed under the action of $N$ independent super-Pauli strings - the so-called ‘super-stabilizers’. In our state-like notation we have $N$ super-stabilisers $\mathbf{O}_{\alpha}$ such that $\mathbf{O}_{\alpha}{\bf[W\rangle}={\bf[W\rangle}$. The super-stabilisers of ${\bf[W\rangle}$ form a $2^{N}$ dimensional group under multiplication, and by independent we mean that they form a choice of generating set for the super-stabiliser group, and none can be generated from the others. A stabiliser operator is uniquely defined by its super-stabiliser group, and vice versa.
@@ III Fast Scrambling of Operators
The previous work Blake and Linden (2020) performed initial studies of operator entanglement in super-Clifford circuits, and in particular demonstrated that super-Clifford circuits were capable of generating a large amount of operator entanglement starting from an initial product operator $W(0)=X_{1}\dots X_{N}$. A central question in the context of scrambling, that was not discussed in detail in Blake and Linden (2020), is to understand how quickly operator entanglement can be generated by super-Clifford circuits.
A key motivation for this question is provided by the fast scrambling conjecture Sekino and Susskind (2008), which hypothesised a lower bound on the ‘scrambling time’ $t_{*}$ of many-body quantum systems - that is the time taken for the system to scramble initially local quantum information across many degrees of freedom. Specifically, in a system of $N$ qubits, Sekino and Susskind (2008) introduced the concept of a ‘fast scrambler’, i.e. a system which scrambles in a time
@@ III.1 Scrambling time from operator entanglement
with $\epsilon>0$ a fixed constant independent of $N$. This definition of scrambling time is entirely analogous to that used in Bentsen et al. (2019) for state entanglement.
@@ III.3 Operator entanglement entropy
Figure 2: The near saturation behaviour of the operator entanglement entropy (averaged over 500 iterations) for a parallel-processing super-Clifford circuit for a range of system sizes $N$. For $N\geq 1560$ we find $\mathrm{ln}(\Delta S_{1/4}(t))$ is well described by a straight line independent of $N$. This is consistent with the scaling form (18) implying the system is fast scrambling.
@@ IV.3 Numerical results for OTOCs
Figure 4: The left hand panel shows OTOCs (22) for $V(0)=C_{3}$ in a parallel processing super-Clifford circuit with $N=120$ qubits, averaged over $1000$ realisations of the circuit. The plots are shown for $W(0)=D_{1}X_{2}\dots X_{N}$ with $D_{1}=X_{1}$ (mauve) and $D_{1}=Y_{1}$ (green). The right hand panel shows the fraction of realisations of the circuit for which $F(t)$ is not given by the scrambled value $F(t)=1/2$ (out of $2000$ realisations with $D_{1}=X_{1}$).
@@
Figure 5: The left hand panel shows the OTOC averaged over $250$ realisations of a parallel processing super-Clifford circuit with $N=1000$ qubits, for operators $V(0)=C_{3}T_{3}$ and $W(0)=D_{1}\dots X_{N}$ with $D_{1}=X_{1}$ (mauve) and $D_{1}=Y_{1}$ (green). The right hand panel shows the OTOC, averaged over $250$ realisations, for $V(0)=C_{3}$, $D_{1}=X_{1}$ for $N=(120,240,360,480)$ qubits.
@@ V Discussion
An important motivation for future work is to understand if the results of Blake and Linden (2020) and this paper can lead to insights into other fast scrambling systems with a ‘large-$N$’ limit, such as holographic quantum field theories or the SYK model. As we have emphasised in the main text, the operators whose scrambling we can currently study using the super-Clifford formalism are ‘many-body operators’ with non-trivial support on each qubits - specifically linear combinations of Pauli strings with $X$ or $Y$ at each site. The majority of studies of scrambling in SYK models and holography so far (e.g. Shenker and Stanford (2014); Roberts et al. (2015, 2018); Maldacena and Stanford (2016); Kitaev and Suh (2018)) have taken place for a qualitatively distinct class of operators - few-body operators with non-trivial support only on a small number of qubits (and identities elsewhere).
@@ Acknowledgements.
AT acknowledges support from UK Engineering and Physical Sciences Research Council (EP/SO23607/1). MB acknowledges support from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe guarantee (EP/Y00468X/1). NL gratefully acknowledges support from the UK Engineering and Physical Sciences Research Council through grants EP/R043957/1, EP/S005021/1, EP/T001062/1. The data and code that supports the findings of this article are openly available at Blake et al. (2024).

### v2

@@ II Super-Clifford circuits
In the context of our super-Clifford circuits, the familiar notion of a stabiliser state is replaced by a ‘stabiliser operator’. Specifically, for an $N$ qubit system, we define a stabiliser operator as an operator in $\mathcal{S}$ which is fixed under the action of $N$ independent super-Pauli strings - the so-called ‘super-stabilizers’. In our state-like notation we have $N$ super-stabilisers $\mathbf{O}_{\alpha}$ such that $\mathbf{O}_{\alpha}{\bf[W\rangle}={\bf[W\rangle}$. The super-stabilisers of ${\bf[W\rangle}$ form a $2^{N}$ dimensional group under multiplication, and by independent we mean that they form a choice of generating set for the super-stabiliser group, and none can be generated from the others. A stabiliser operator is uniquely defined by its super-stabiliser group, and vice versa.
@@ III Fast Scrambling of Operators
The previous work Blake and Linden (2020) performed initial studies of operator entanglement in super-Clifford circuits, and in particular demonstrated that super-Clifford circuits were capable of generating a large amount of operator entanglement starting from an initial product operator $W(0)=X_{1}\dots X_{N}$. We note that this implies that the physics of such circuits is qualitatively distinct from conventional Clifford circuits. In particular whilst Clifford circuits are capable on generating state entanglement from an initial product state, they are not capable of generating operator entanglement starting from a single Pauli string.
A central question in the context of scrambling, that was not discussed in detail in Blake and Linden (2020), is to understand how quickly operator entanglement can be generated by super-Clifford circuits. A key motivation for this question is provided by the fast scrambling conjecture Sekino and Susskind (2008), which hypothesised a lower bound on the ‘scrambling time’ $t_{*}$ of many-body quantum systems - that is the time taken for the system to scramble initially local quantum information across many degrees of freedom. Specifically, in a system of $N$ qubits, Sekino and Susskind (2008) introduced the concept of a ‘fast scrambler’, i.e. a system which scrambles in a time
@@ III.1 Scrambling time from operator entanglement
with $\epsilon>0$ a fixed constant independent of $N$. This definition of scrambling time is entirely analogous to that used in Bentsen et al. (2019) for state entanglement [^2] We note that it is also common to define a scrambling time in terms of OTOCs of local operators, e.g. as in  Keselman et al. (2021); Zhou and Chen (2019). We are not aware of a direct relationship between these alternative definitions and those based on entanglement entropy as used here or in Bentsen et al. (2019)..
@@ III.3 Operator entanglement entropy
Figure 2: (Left) $\mathrm{ln}(\Delta(S_{1/4}(t)))$ for a range of system sizes $N\geq 720$ averaged over $500$ iterations of a parallel-processing super-Clifford circuit. At an N independent timescale (shown by the red dotted line) the data is approximately described by a straight line with N independent parameters. (Right) The same data but only for times $t\geq 50$, including a straight line fit (black line) on the data obtained for $N=3000$. The straight line with independent parameters observed here is consistent with the scaling form (18) implying the system is fast scrambling.
@@
To conclude this section, it is instructive to clarify the origin of the fast scrambling behaviour $t_{*}\sim\ln(N)$ in the operator entanglement of these circuits. It was argued in Bentsen et al. (2019), in the context of state entanglement, that a fast scrambling bound on state entanglement arises due to a slow-down in the rate of entropy production at late times. Specificially, it was argued that the entropy of macroscopic subregions is governed by (at best) a slow exponential decay towards maximal entanglement. The numerical results in this paper, specifically the fit to the functional form (18) can be viewed as an explicit realisation of this behaviour in the context of the (averaged) operator entanglement of this parallel-processing super-Clifford circuits.
@@ IV.3 Numerical results for OTOCs
Figure 4: The left hand panel shows OTOCs (22) for $V(0)=C_{3}$ in a parallel processing super-Clifford circuit with $N=120$ qubits, averaged over $1000$ realisations of the circuit. The plots are shown for $W(0)=D_{1}X_{2}\dots X_{N}$ with $D_{1}=X_{1}$ (mauve dashed line) and $D_{1}=Y_{1}$ (yellow solid line). The right hand panel shows the fraction of realisations of the circuit for which $F(t)$ is not given by the scrambled value $F(t)=1/2$ (out of $2000$ realisations with $D_{1}=X_{1}$).
@@
Figure 5: The left hand panel shows the OTOC averaged over $250$ realisations of a parallel processing super-Clifford circuit with $N=1000$ qubits, for operators $V(0)=C_{3}T_{3}$ and $W(0)=D_{1}\dots X_{N}$ with $D_{1}=X_{1}$ (mauve dashed line) and $D_{1}=Y_{1}$ (yellow solid line). The right hand panel shows the OTOC, averaged over $250$ realisations, for $V(0)=C_{3}$, $D_{1}=X_{1}$ for $N=(120,240,360,480)$ qubits.
@@ V Discussion
Throughout this paper we have been concerned with the scrambling of ‘many-body’ operators $W(t)$ which have global support on our system of $N$ qubits. Whilst the majority of studies in the literature focus on scrambling of local few-body operators, the scrambling of similar global operators has previously been studied in Swingle et al. (2016). In particular, Swingle et al. (2016) presented results for studies of OTOCs where both $V(0)$ and $W(0)$ can be viewed as a global product of rotation operators in an ensemble of spins. [^4] Other studies of operators with global support include Lin and Motrunich (2018); Zhou and Swingle (2023); Kukuljan et al. (2017). The global operators studied in Zhou and Swingle (2023) and Kukuljan et al. (2017) are qualitatively different to those studied in this paper, as they can be expressed as sums of local operators. In this paper we studied a slightly different form of the OTOC, where $W(0)$ was a global operator and $V(0)$ a local operator. Surprisingly we found that such OTOCs between many-body and local operators scrambled rapidly within an $\mathcal{O}(1)$ time in our parallel-processing super-Clifford circuit. It would be interesting to understand if this behaviour is generic for such OTOCs in fast scrambling systems. A further interesting direction is to use super-Clifford circuits to study more general types of OTOCs involving many-body operators. The method we have introduced to compute OTOCs in super-Clifford circuits can also be applied to OTOCs where both $W(0)$ and $V(0)$ are operators with global support, as in Swingle et al. (2016), and studying this is an interesting avenue for future work.
An important motivation for future work is to understand if the results of Blake and Linden (2020) and this paper can lead to insights into other fast scrambling systems with a ‘large-$N$’ limit, such as holographic quantum field theories or the SYK model. As we have emphasised, the operators whose scrambling we can currently study using the super-Clifford formalism are ‘many-body operators’ with non-trivial support on each qubits - specifically linear combinations of Pauli strings with $X$ or $Y$ at each site. The majority of studies of scrambling in SYK models and holography so far (e.g. Shenker and Stanford (2014); Roberts et al. (2015, 2018); Maldacena and Stanford (2016); Kitaev and Suh (2018)) have taken place for a qualitatively distinct class of operators - few-body operators with non-trivial support only on a small number of qubits (and identities elsewhere).
@@ Acknowledgements.
AT acknowledges support from UK Engineering and Physical Sciences Research Council (EP/SO23607/1). MB acknowledges support from UK Research and Innovation (UKRI) under the UK government’s Horizon Europe guarantee (EP/Y00468X/1). NL gratefully acknowledges support from the UK Engineering and Physical Sciences Research Council through grants EP/R043957/1, EP/S005021/1, EP/T001062/1. The data and code that supports the findings of this article are openly available at Blake et al. (2025). For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising from this submission.